# Provable Low-Frequency Bias of In-Context Learning of Representations

### v1

@@ Appendix E Integrating Other Transformer Components
In this section, we discuss how our theoretical results can extend to more complex and realistic Transformer architectures beyond the simplified model described in Algorithm 1. We first note that in our framework, any neuron-wise transformation (i.e., operations that apply independently to the representation of each token) can be absorbed into the definition of the great mapping $\sigma$. This includes FFNs as well as normalizations such as LayerNorm or RMSNorm. Therefore, here we focus here two architectural components not yet discussed: residual connections and multi-head attentions.
@@
The most critical step in the proof of Theorem 2 is Lemma 8, which establishes that applying a nice attention map to a good sequence results in another good sequence, and that the attention operates implicitly on the latent representations to which the sequence converges. Introducing residual connections here is straightforward: with residual connection, eq. 33 would become
@@
which is simply adding $1$ to the $\rho_{A}^{(\ell)}$ coefficient. This modification only affects the $\delta_{q}$ term in Theorem 2, which becomes
@@
Lemma 7 shows that any Lipschitz combination of good sequences remains a good sequence. Since multi-head attention can be viewed as a Lipschitz combination of multiple single-head attentions, it follows that a multi-head attention mechanism also satisfies Lemma 8, as long as each individual head does. All other parts of the theoretical framework extend accordingly. Notice that the coefficients in Lemma 8 may differ by constant factors in the multi-head case, but this does not affect the asymptotic conclusions in Theorem 2, which only concern limiting behavior.
@@ Appendix F Empirical Verification
As discussed in the main text, Theorem 2 relies on a relatively strong structural assumption about the attention maps. It is therefore essential to verify whether these assumptions hold in practice. In this section, we empirically examine this question.
Specifically, in Figure 5, we compute the proportion of A,B,T type attention connections defined in Section 5. Note that Type O connections (i.e., connecting to arbitrary tokens) are excluded from this analysis, as they cover all positions. For each head, we compute the fraction of attention weights (across all layers) that fall into types A, B, or T. Overlapping cases (e.g., the connection from the second token in the sequence to the first one can be considered both as B and as T type) are counted only once. The figure shows that a large proportion of attention weights (> $72\%$ in total) indeed falls into these structured types, lending empirical support to our theoretical assumptions.

### v2

@@ Appendix E Integrating Other Transformer Components
In this section, we discuss how our theoretical results can extend to more complex and realistic Transformer architectures beyond the simplified model described in Algorithm˜1. We first note that in our framework, any neuron-wise transformation (i.e., operations that apply independently to the representation of each token) can be absorbed into the definition of the great mapping $\sigma$. This includes FFNs as well as normalizations such as LayerNorm or RMSNorm. Therefore, here we focus here two architectural components not yet discussed: residual connections and multi-head attentions.
@@
The most critical step in the proof of Theorem˜2 is Lemma˜8, which establishes that applying a nice attention map to a good sequence results in another good sequence, and that the attention operates implicitly on the latent representations to which the sequence converges. Introducing residual connections here is straightforward: with residual connection, eq.˜33 would become
@@
which is simply adding 111 to the $\rho_{A}^{(\ell)}$ coefficient. This modification only affects the $\delta_{q}$ term in Theorem˜2, which becomes
@@
Lemma˜7 shows that any Lipschitz combination of good sequences remains a good sequence. Since multi-head attention can be viewed as a Lipschitz combination of multiple single-head attentions, it follows that a multi-head attention mechanism also satisfies Lemma˜8, as long as each individual head does. All other parts of the theoretical framework extend accordingly. Notice that the coefficients in Lemma˜8 may differ by constant factors in the multi-head case, but this does not affect the asymptotic conclusions in Theorem˜2, which only concern limiting behavior.
@@ Appendix F Empirical Verification
As discussed in the main text, Theorem˜2 relies on a relatively strong structural assumption about the attention maps. It is therefore essential to verify whether these assumptions hold in practice. In this section, we empirically examine this question.
[…]